I have a 50% solution of hydrogen peroxide and I want to make 1000mls of 30% hydrogen peroxide. How much water and concentrate should I use?
Found 2 solutions by mananth, jorel1380:
Answer by mananth(16946)

percent percentage ---------------- quantity
Hydrogen peroxide 50 ---------------- x ml
water 0 ---------------- 1000 - x ml
Mixture 30.00% ---------------- 1000

50 x + 0 ( 1000 - x ) = 30 * 1000

50 x + 0 - 0 x = 30000
50 x - 0 x = 30000 - 0
50 x = 30000
/ 50
x= 600ml 50.00% Hydrogen peroxide
400 ml water

Answer by jorel1380(3719)

1000 mls x 30%=300 mls
1000 mls - 300 mls = 700 mls water
.5(x)+1(1000-x)=700
.5x+1000-x=700
.5x=300
x=300/.5=600
You need 600 ml of .5 solution and 400 ml of pure water..